OpenStudy (anonymous):

solve the equation. x^3-x^2-3x+3=0

jimthompson5910 (jim_thompson5910):

x^3-x^2-3x+3=0 x^2(x-1)-3(x-1)=0 (x^2-3)(x-1) = 0 x^2-3 = 0 or x-1 = 0 x^2 = 3 or x = 1 x = sqrt(3), x = -sqrt(3) or x = 1